On Symmetric Block Designs (81, 16, 3) with Involutory Automorphisms Fixing 17 Points

Ljubo Marangunić, Zagreb, Croatia

Keywords: block design, fixed point, fixed structure, orbit, orbital structure AMS Subject classification: 05B05 Abstract: We prove the nonexistence of the (81, 16, 3)-symmetric block design admitting an involutory automorphism fixing 17 points.